# A Wasserstein perspective of Vanilla GANs

### v1

@@ 6 Wasserstein GAN
The same analysis can be applied to Wasserstein-type GANs. The constrained on the Hölder constant can be weakened, as we do not need Theorem 3.2. Note that this does not impact the rate, but the constant. Define the Wasserstein-type GAN with discriminator class $\mathcal{W}$ as
@@
For $0<\alpha<1,$ $\Gamma>2d^{1/p},M>0$ and $d>2\alpha$ and $n>2^{\frac{d}{\alpha}}$ choose $\mathcal{G}$ as the set of ReLU networks with at most $c\cdot(\log(n)+1)$ layers, $c\cdot n^{{\alpha}}(\log(n)+1)$ neurons an nonzero weights and $\mathcal{W}^{\prime}$ as the set of ReLU networks with at most $c\cdot\log(n)$ layers and $c\cdot n^{\frac{\alpha}{(1-\alpha)}}\log^{2}(n)$ nonzero weights and neurons, where $c$ is a constant depending on $d,d^{*},\Gamma$ and $\alpha$ The empirical risk minimizer with $\mathcal{W}=\mathcal{W}^{\prime}\cap\mathcal{H}^{\alpha}(\Gamma)$
@@
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]\leq\min\Big{\{}c\cdot n^{-\frac{\alpha}{d}}+\inf_{G^{*}\mathrm{mb.}}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)}),\;c\cdot n^{-\frac{\alpha}{d^{*}}}+\inf_{G^{*}\in\operatorname{Lip}(M,\mathcal{Z})}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})\Big{\}}.$
@@
Compared to Corollary 5.4 the rate improves to $n^{-\alpha/d^{*}}$ for any $\alpha<1$. This upper bound thus coincides with the lower bound in Tang & Yang, (2022, Theorem 1) up to an arbitrary small polynomial factor. Our rate does not depend exponentially on the number of layers like the results of Liang, (2017), Huang et al., (2022) and we use non-smooth simple ReLU networks compared to smooth ReQU networks in Stéphanovitch et al., (2023) or group sort networks in Biau et al., (2021). In addition, are able to evaluate in the Wasserstein metric rather than the $\mathcal{H}^{\alpha}$ metric.
7 Proofs
7.1 Proof for Section 2
@@ 7.2 Proofs for Section 3
$\leq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>\log(2)\end{subarray}}\mathbb{E}[\psi(f(W(X)))+\psi(-f(W(Y)))]$
$=\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>\log(2)\end{subarray}}\mathbb{E}[\psi(f(W(X)))-W(Y)].$
@@
$\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})\leq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>\log(2)\end{subarray}}\mathbb{E}[W(X)-W(Y)]-\frac{e^{B}}{(2e^{B}-1)^{2}}\mathbb{E}[W(X)^{2}].\qed$
@@
Let $W^{*}\in\underset{\begin{subarray}{c}W\in\operatorname*{Lip}(1,\log(2-2/L))\end{subarray}}{\arg\max}\mathbb{E}[W(X)-W(Y)].$ Then $\delta W^{*}\in\operatorname*{Lip}(1,\log(2-2/L))$ for all $\delta\in(0,1]$ and we can conclude
@@
$\geq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(\log(2-2/L))d^{-1/p},\infty)\end{subarray}}\mathbb{E}[W(X)-W(Y)]=\frac{\log(2-2/L))}{d^{1/p}}\mathsf{W}_{1}(\mathbb{P},\mathbb{Q}),$
we conclude the claimed lower bound for
@@
For the upper bound we use Lemma 7.1 with $\mathcal{W}=\operatorname*{Lip}(L).$ Since for $W\in\operatorname*{Lip}(L)$ the function $-\log\big{(}1+e^{-W(\cdot)}\big{)}\in\operatorname*{Lip}(L)$ we conclude
@@ 7.3 Proofs for Section 4
Since $G^{*}$ was arbitrary, Theorem 4.1 yields for some constant $c$
@@
Using Jensen’s inequality, we can bound the stochastic error term by
@@
Note that $\mathbb{E}[|X_{i}-G^{*}(Z_{i})|_{p}]=\mathsf{W}_{1}(\mathbb{P}^{G^{*}(Z)},\mathbb{P}^{*})$ by the duality formula of $\mathsf{W}_{1}$ used in this work, see Villani, (2008, Definition 6.2 and Remark 6.5). For $\mathbb{E}[W_{1}(\mathbb{U}_{n},\mathbb{U})]$, we can exploit the convergence rate for the empirical distribution as in Corollary 4.2, but now in the $d^{*}$ dimensional latent space $\mathcal{Z}$. Therefore, there exists a $c^{\prime}$ such that
@@
7.4 Proofs of Theorem 5.1 and Theorem 5.2
@@
due to Lipschitz continuity of $x\mapsto-\log((1+e^{x})/2)$. From (22) we deduce for $\mathcal{W}\subset\operatorname*{Lip}(L,B)$
@@
Since $\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})$ can be estimated as in Theorem 4.3, we only need to bound $\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})$. For $L=\Gamma,B=\Gamma/2$ we have $\operatorname*{Lip}(L,B)\subset\mathcal{H}^{\alpha}(\Gamma)$, $\alpha\in(0,1)$, and the assumptions of Theorem 3.2 are satisfied. Therefore for every $\alpha\in(0,1)$
@@
To bound the first term, we apply Lemma 7.1 and $\{-\log(1+e^{-W(\cdot)})\mid W\in\mathcal{H}^{\alpha}(\Gamma)\}\subset\mathcal{H}^{\alpha}(\Gamma)$ to obtain
@@
$\leq\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}_{\begin{subarray}{c}\hat{X}\sim\mathbb{P}_{n}\\ \end{subarray}}[W(\hat{X})-W(G(Z))]$
@@
Theorem 5.3 is very close to Theorem 4.1 by Gühring et al., (2020), which however applies only to functions $f$ which are at least twice (weakly) differentiable. Our proof can thus build on numerous auxiliary results and arguments from Gühring et al., (2020). We basically keep the proof structure of Gühring et al., (2020) which in turn relies on Yarotsky, (2017).
@@
$\Big{\|}f-\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}c_{f,m}\Big{\|}_{W^{k,\infty}((0,1)^{d})}$ $\leq C_{6}\Big{(}\frac{1}{N}\Big{)}^{(1-k)}\sup_{\tilde{m}\in\{0,\ldots,N\}^{d}}\Big{(}\sup_{\begin{subarray}{c}m\in\{0,\ldots,N\}^{d},\\ |m-\widetilde{m}|_{\infty}\leq 1\end{subarray}}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m,N})}\Big{)}$ (29)
@@
Lemma 7.4.
For any $\varepsilon\in(0,1/2)$ there is a neural network $\Phi_{\varepsilon}$ with ReLU activation function such that for $(c_{f,m})_{m}$ from Lemma 7.3, there is a constant $C>0$ such that
@@
the number of layers is at most $\lceil C_{1}\log_{2}(\varepsilon^{-1})\rceil,$ the number of nonzero weights is at most $\lceil C_{2}(N+1)^{d}\log_{2}^{2}(\varepsilon^{-1})\rceil$ and the number of neurons is at most $\lceil C_{3}(N+1)^{d}(\log_{2}^{2}(\varepsilon^{-1})\vee\log_{2}(\varepsilon^{-1}))\rceil,$ with $C_{1},C_{2}$ and $C_{3}$ from Lemma 7.2.
@@
From Lemma 7.2 we know that there are neural networks $\Psi_{\varepsilon,m}$ with at most $\lceil C_{1}\log_{2}(\varepsilon^{-1})\rceil$ layers, $\lceil C_{2}(N+1)^{d}\log_{2}^{2}(\varepsilon^{-1})\rceil$ nonzero weights and $\lceil C_{3}(N+1)^{d}(\log_{2}^{2}(\varepsilon^{-1})\vee\log_{2}(\varepsilon^{-1}))\rceil$ neurons that approximate $\phi_{m}$ such that for $k\in\{0,1\}$
@@
The second to last inequality follows from the fact that on $\Omega_{\widetilde{m},N}$ is within the support of $\phi_{m}$ only for $|m-\widetilde{m}|_{\infty}\leq 1.$ The last inequality follows from (31) and the continuity of the extension operator, see (Stein,, 1970, Theorem 5). As in Step 4 of Lemma 7.3, we conclude using Lunardi, (2018, Theorem 1.6)
@@
Combining Lemma 7.3 and Lemma 7.4 with $\|f\|_{W^{1,\infty}}\leq B$ yields for a constant $C>0$ for any $\tilde{\varepsilon}\in(0,1/2)$ that
@@
where $\tilde{\varepsilon}$ determines the approximation accuracy in Lemma 7.4. For
@@
From Lemma 7.2 we know that there is a ReLU network with no more than $1+\lceil C_{1}\log_{2}(\tilde{\varepsilon}^{-1})\rceil$ layers, $N^{d}+\lceil C_{2}(N+1)^{d}\log_{2}^{2}(\tilde{\varepsilon}^{-1})\rceil$ nonzero weights and $1+\lceil C_{3}(N+1)^{d}(\log_{2}^{2}(\tilde{\varepsilon}^{-1})\vee\log_{2}(\tilde{\varepsilon}^{-1}))\rceil$ neurons with the required properties. Inserting (33) and assuming $CB>\frac{1}{2}$ yields
@@
We get with the same reasoning as in the proof of (32)

### v2

@@ 6 Wasserstein GAN
The same analysis can be applied to Wasserstein-type GANs. The constrained on the Hölder constant can be weakened, as we do not need Theorem 3.2. Note that this does not impact the rate, but the constant. Define the Wasserstein-type distance with discriminator class $\mathcal{W}$ as
@@
For $0<\alpha<1,$ $\Gamma>1,M>0$ and $d>2\alpha$ and $n>2^{\frac{d}{\alpha}}$ choose $\mathcal{G}$ as the set of ReLU networks with at most $\lceil c\cdot\log(n)\rceil$ layers, $\lceil c\cdot n\log(n)\rceil$ nonzero weights and $\lceil c\cdot n\log(n)\rceil$ neurons and $\mathcal{W}^{\prime}$ as the set of ReLU networks with at most $\lceil c\cdot\log(n)\rceil$ layers, $\lceil c\cdot n^{\frac{\alpha}{(1-\alpha)}}\log^{2}(n)\rceil$ nonzero weights and $\lceil c\cdot n^{\frac{\alpha}{(1-\alpha)}}\log^{2}(n)\rceil$ neurons, where $c$ is a constant depending on $d,d^{*},\Gamma,M$ and $\alpha$. The empirical risk minimizer with $\mathcal{W}=\mathcal{W}^{\prime}\cap\mathcal{H}^{\alpha}(\Gamma)$
@@
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]\leq c\cdot n^{-\frac{\alpha}{d^{*}}}+\inf_{G^{*}\in\operatorname{Lip}(M,\mathcal{Z})}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)}).$
Compared to Corollary 5.4 the rate improves to $n^{-\alpha/d^{*}}$ for any $\alpha<1.$ The number of observations necessary for the theorem to hold, the size of the discriminator network and the lower bound for $\Gamma$ decrease. Note that $\Gamma$ does not effect the rate, but the constants. In case there exists a $G^{*}$ such that $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})=0,$ this upper bound coincides with the lower bound in Tang & Yang, (2023, Theorem 1) up to an arbitrary small polynomial factor.
@@
Our rate does not depend exponentially on the number of layers like the results of Liang, (2017), Huang et al., (2022) and we use non-smooth simple ReLU networks compared to smooth ReQU networks in Stéphanovitch et al., (2023) or group sort networks in Biau et al., (2021).
7 Numerical illustration
The results in Section 5 and Section 6 were obtained under the assumption that the discriminator class consists of Lipschitz networks. In the context of image generation, these findings align with the results of Zhou et al., (2019), Miyato et al., (2018), Kodali et al., (2017), and Fedus et al., (2017). Furthermore, Fedus et al., (2017) demonstrated in a two-dimensional experiment that a Vanilla GAN with a gradient penalty (and, consequently, a lower Lipschitz constant) can be effective in scenarios where the measures $\mathbb{P}^{*}$ and $\mathbb{P}^{\hat{G}(Z)}$ are singular.
This section presents a transparent and accessible example that confirms our theoretical findings and especially demonstrates how imposing a Lipschitz constant on the discriminator stabilizes the Vanilla GAN. Additionally, it demonstrates the capacity of the Vanilla GAN to detect a lower dimensional manifold. In order to monitor rates of convergence, it is necessary to at least approximately evaluate $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})$. Therefore, we study the numerical performance of the Vanilla GAN in a simulation setting, where the true data distribution is known by construction.
In this work, the Wasserstein distance is employed as the metric for measuring the rate of convergence. In practice, the Wasserstein distance is only computable in the one-dimensional case. To investigate multivariate distributions, we approximated the Wasserstein distance by averaging the Wasserstein distance on the marginals.
In order to model the distribution $\mathbb{P}^{*}$ of a lower dimensional manifold, we employed a one-dimensional uniform distribution on the graph of the function $x\mapsto\sin(4\pi x)$ on the diagonal of the two-dimensional unit cube, resulting in a three-dimensional distribution. For the latent distribution, we used the one-dimensional normal distribution. Consequently, the dimensions of the lower dimensional manifold and the latent space are identical.
For the discriminator, we used a neural network with four layers of width $128$ concatenated to a sigmoid function. For the generator, we used a neural network with three layers of width $64$ In order to preserve as much alignment as possible with the theoretical result, we used plain ReLU activations. Each training consisted of $30000$ training iterations. We used the Adam optimizer (Kingma & Ba,, 2014) with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and a learning rate of $\gamma=0.0005$. When the number of observations exceeded $512,$ we used minibatches of that size in each iteration. We updated generator and discriminator alternating.
Three snapshots of the training of the Vanilla GAN for samples sizes $n=100$ and $n=1000$ are given in Figure 1, respectively. The difference between Figure 1(c) and Figure 1(f) is solely due to the number of observations. The observations in Figure 1(a) to Figure 1(c) cover the manifold to a lesser extent than the observations in Figure 1(d) to Figure 1(f). This corresponds to a larger stochastic error.
(a) Step $0,$ $100$ obs.
(b) Step $15000,$ $100$ obs.
(c) Step $30000,$ $100$ obs.
(d) Step $0,$ $1000$ obs.
(e) Step $15000,$ $1000$ obs.
(f) Step $30000,$ $1000$ obs.
Figure 1: Training of Vanilla GAN with weight clip using $100$ observations (first row) or $1000$ observations (second row). Red dots show $1000$ generated samples, green dots show the observations used for the training. The blue line is the one-dimensional manifold.
To maintain the Lipschitz constant within a controllable range, we implemented the simple weight clipping mechanism of Arjovsky et al., (2017), limiting each weight to a value of $0.5$. It is important to note that the network used in the unclipped case is also Lipschitz continuous, however, we do not have control over this Lipschitz constant. Given the width and depth parameters used in this study, it is evident that the Lipschitz constant of the clipped network remains relatively high and is considerably distinct from the theoretical value typically employed in Wasserstein GANs. However, a smaller Lipschitz constant requires an adjustment to the learning rate. Otherwise the weights are likely to remain at their maximum absolute value. This affects the experiment in several other ways. To ensure a fair and accurate comparison between the clipped and unclipped scenarios, we kept the learning rate consistent.
The results are summarized in Figure 2.
Figure 2: Marginal $\mathsf{W}_{1}$ distance depending on number of observations. Thick line shows the average over $50$ independent runs, ribbons show the first to third quartile.
As predicted by our theory, the averaged marginal Wasserstein distance between the generated distribution and the true data distribution decays approximately as $n^{-1/2}$ for $n\in\{10,100,1000\}$. While we see a clear improvement with $10000$ observations, the additional gain is limited by the optimization error, since the manifold is already densely covered for $1000$ observations.
It is apparent that controlling the Lipschitz constant overall stabilizes the training process, resulting in less variability in the results. In certain cases, the GAN without weight clipping can achieve the same level of effectiveness. This does not negate the outcome. Since the discriminator without clipped weights is still Lipschitz continuous (with a large Lipschitz constant), the theoretical limitations of Vanilla GANs without restricted discriminator classes do not directly translate to practice. This, combined with the finite nature of the implementations, ultimately resulted in the empirical success of these models. The variability between different simulation runs is described by the first to the third quartile in Figure 2 which again confirms a more stable behavior of the clipped algorithm.
Figure 3: Percentage of generated samples with euclidean distance to manifold greater than $\varepsilon$ using $1000$ observations and a discriminator with $0.5$ clip. Transparent lines show the individual runs, thick line shows the average over $50$ runs.
Figure 3 demonstrates the high degree of precision with which the generated samples concentrate on the low-dimensional support of the true data distribution. Our experiments show that this concentration holds true across all sample sizes and can be observed in both the clipped and unclipped case. However, a high concentration does not necessarily indicate that the generated distribution is an accurate imitation of the unknown distribution with respect to the Wasserstein distance. Consequently, Figure 3 is only informative in conjunction with Figure 2.
Additionally, we investigated the use of a space $\mathbb{U}$ of the same dimension as the ambient space. Our observations indicated that the Vanilla GAN is still capable of identifying the lower dimensional subspace with reasonable efficacy.
8 Discussion and limitations
Our analysis demonstrates that GANs originally build on too sensitive distribution distances such as the Jensen-Shannon distance can be improved via a Lipschitz constraint in the discriminator class. This insight might be also applicable to other GANs, e.g. $f$-GANs by Nowozin et al., (2016), which rely on a divgerence which cannot disciminate between different singular distributions and is thus not suitable for a dimension reduction setting. Overall we conclude that the choice of the discriminator class is much more important for the data generation capabilities than choice of the loss function which is typically dictated by some distance. Moreover, our analysis of the approximation error of the discriminator is not limited to the Vanilla GANs, but also applicable to optimal transport based GANs as demonstrated for the Wasserstein GAN.
There are several potential avenues for further development of the results presented in this paper. This includes especially the mentioned potential implications to other types of GANs. While our analysis was limited to ReLU feedforward networks, one advancement of research on neural networks is the use of more sophisticated networks whose statistical analysis is not yet settled. In context of Wasserstein GANs, see for example Radford et al., (2015).
Furthermore, inclusion of a bound on the Lipschitz constant (and not only the Hölder constant) would allows for direct application of Theorem 5.1, thereby eliminating the need to include the parameter $\alpha$ and thus improve the rates. Additionally it would be interesting, whether there are conditions that allow for a faster rate of convergence for the Vanilla GAN in some cases (excluding scenarios as in Example 3.3).
The experiments also demonstrated that the GAN is capable of detecting data from a lower dimensional manifold if the latent space is of the same dimension as the ambient space. The proof of Theorem 4.3 is contingent upon the dimension of the latent space. If the dimension of the latent space is chosen to be too small, then $\inf_{G^{*}\in\operatorname{Lip}(M,\mathcal{Z})}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})$ will be large. If the dimension of the latent space is chosen too large, this does not result in a deterioration of $\inf_{G^{*}\in\operatorname{Lip}(M,\mathcal{Z})}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})$ compared to the optimal choice. However, the rate that was proven then still depends on the higher latent dimension. Therefore rates that are adaptive to the unknown intrinsic dimension, potentially benefiting from results like Berenfeld & Hoffmann, (2021), would be interesting.
Bounds in other distances suitable to the dimension reduction setting are also of high interest. For example, the Wasserstein-$2$ metric is slightly stronger than the Wasserstein-$1$ metric (in the sense that $\mathsf{W}_{1}\leq\mathsf{W}_{2}$, see (Villani,, 2008, Remark 6.6)). Our proofs rely on the duality of the Wasserstein-$1$ distance, hence they cannot be translated directly to the Wasserstein-$2$ distance.
Finally, the objective of this study was to examine statistical perspectives, and thus, the optimization problem was not addressed. In the proofs, we employ the global minimizer and maximizer. Since we face a non-convex optimization problem gradient based methods may suffer from a considerable optimization error, especially for high-dimensional parameter spaces. Incorporating this optimization error would align more closely with real-world scenarios.
Appendix A Appendix
A.1 Proof for Section 2
@@ A.2 Proofs for Section 3
$\leq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>-\log(2)\end{subarray}}\mathbb{E}[\psi(f(W(X)))+\psi(-f(W(Y)))]$
$=\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>-\log(2)\end{subarray}}\mathbb{E}[\psi(f(W(X)))-W(Y)].$
@@
$\mathsf{V}_{\operatorname*{Lip}(L,B)}(\mathbb{P},\mathbb{Q})\leq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(L,B)\\ W(\cdot)>-\log(2)\end{subarray}}\mathbb{E}[W(X)-W(Y)]-\frac{e^{B}}{(2e^{B}-1)^{2}}\mathbb{E}[W(X)^{2}].\qed$
@@
Let $W^{*}\in\underset{\begin{subarray}{c}W\in\operatorname*{Lip}(1,\log(2-2/L))\end{subarray}}{\arg\max}\mathbb{E}[W(X)-W(Y)].$ This element exists by (Villani,, 2008, Theorem 5.10 (iii)). Then $\delta W^{*}\in\operatorname*{Lip}(1,\log(2-2/L))$ for all $\delta\in(0,1]$ and we can conclude
@@
$\geq\sup_{\begin{subarray}{c}W\in\operatorname*{Lip}(\log(2-2/L)d^{-1/p},\infty)\end{subarray}}\mathbb{E}[W(X)-W(Y)]$
$=\frac{\log(2-2/L))}{d^{1/p}}\mathsf{W}_{1}(\mathbb{P},\mathbb{Q}).$
Hence we can conclude the claimed lower bound for
@@
For the upper bound we use Lemma A.1 with $\mathcal{W}=\operatorname*{Lip}(L).$ Since for $W\in\operatorname*{Lip}(L)$ the function $-\log\big{(}1+e^{-W(\cdot)}\big{)}\in\operatorname*{Lip}(L)$ we conclude
@@ A.3 Proofs for Section 4
Since $G^{*}$ was arbitrary Theorem 4.1 yields for some constant $c$
@@
Here the infimum can be used as we can increase the constant $c$ multiplied to both terms by an arbitrary small $\varepsilon>0$ to account for the possibly infinitesimal smaller value. Using Jensen’s inequality, we can bound the stochastic error term by
@@
Note that $\mathbb{E}[|X_{i}-G^{*}(Z_{i})|_{p}]=\mathsf{W}_{1}(\mathbb{P}^{G^{*}(Z)},\mathbb{P}^{*})$ by the duality formula of $\mathsf{W}_{1}$ used in this work, see Villani, (2008, Definition 6.2 and Remark 6.5). For $\mathbb{E}[W_{1}(\mathbb{U}_{n},\mathbb{U})]$, we can exploit the convergence rate for the empirical distribution as in Corollary 4.2, but now in the $d^{*}$-dimensional latent space $\mathcal{Z}$. Therefore, there exists a $c^{\prime}$ such that
@@
A.4 Proofs of Theorem 5.1 and Theorem 5.2
@@
due to Lipschitz continuity of $x\mapsto-\log((1+e^{x})/2)$.
From (14) we deduce for $\mathcal{W}\subset\operatorname*{Lip}(L,B)$
@@
Since $\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})$ can be estimated as in Theorem 4.3, we only need to bound $\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})$. For $\Gamma>\max(L,2B)$, we have $\operatorname*{Lip}(L,B)\subset\mathcal{H}^{\alpha}(\Gamma)$, $\alpha\in(0,1)$, and the assumptions of Theorem 3.2 are satisfied. Therefore for every $\alpha\in(0,1)$
@@
To bound the first term, we apply Lemma A.1 and $\{-\log(1+e^{-W(\cdot)})\mid W\in\mathcal{H}^{\alpha}(\Gamma)\}\subset\mathcal{H}^{\alpha}(\Gamma)$ to obtain
@@
$\leq\sup_{W\in\mathcal{H}^{\alpha}(\Gamma)}\mathbb{E}_{\begin{subarray}{c}\hat{X}\sim\mathbb{P}_{n}\\ \end{subarray}}[W(\hat{X})-W(G(Z))]$ (15)
@@
Theorem 5.3 is very close to by (Gühring et al.,, 2020, Theorem 4.1), which however applies only to functions $f$ which are at least twice (weakly) differentiable. Our proof can thus build on numerous auxiliary results and arguments from Gühring et al., (2020). We basically keep the proof structure of Gühring et al., (2020) which in turn relies on Yarotsky, (2017).
@@
$\Big{\|}f-\sum_{m\in\{0,\ldots,N\}^{d}}\phi_{m}c_{f,m}\Big{\|}_{W^{k,\infty}((0,1)^{d})}$ $\leq C_{6}\Big{(}\frac{1}{N}\Big{)}^{(1-k)}\sup_{\tilde{m}\in\{0,\ldots,N\}^{d}}\Big{(}\sup_{\begin{subarray}{c}m\in\{0,\ldots,N\}^{d},\\ |m-\widetilde{m}|_{\infty}\leq 1\end{subarray}}\|\tilde{f}\|_{W^{1,\infty}(\Omega_{m,N})}\Big{)}$
@@
Lemma A.4.
For any $\varepsilon\in(0,1/2)$ there is a neural network $\Phi_{\varepsilon}$ with ReLU activation function such that for $(c_{f,m})_{m}$ from Lemma A.3, there is a constant $C>0$ such that
@@
the number of layers is at most $\lceil C_{1}\log_{2}(\varepsilon^{-1})\rceil,$ the number of nonzero weights is at most $\lceil C_{2}(N+1)^{d}\log_{2}^{2}(\varepsilon^{-1})\rceil$ and the number of neurons is at most $\lceil C_{3}(N+1)^{d}(\log_{2}^{2}(\varepsilon^{-1})\vee\log_{2}(\varepsilon^{-1}))\rceil,$ with $C_{1},C_{2}$ and $C_{3}$ from Lemma A.2.
@@
From Lemma A.2 we know that there are neural networks $\Psi_{\varepsilon,m}$ with at most $\lceil C_{1}\log_{2}(\varepsilon^{-1})\rceil$ layers, $\lceil C_{2}(N+1)^{d}\log_{2}^{2}(\varepsilon^{-1})\rceil$ nonzero weights and $\lceil C_{3}(N+1)^{d}(\log_{2}^{2}(\varepsilon^{-1})\vee\log_{2}(\varepsilon^{-1}))\rceil$ neurons that approximate $\phi_{m}$ such that for $k\in\{0,1\}$
@@
The second to last inequality follows from the fact that on $\Omega_{\widetilde{m},N}$ is within the support of $\phi_{m}$ only for $|m-\widetilde{m}|_{\infty}\leq 1.$ The last inequality follows from (21) and the continuity of the extension operator, see (Stein,, 1970, Theorem 5). As in Step 4 of Lemma A.3, we conclude using Lunardi, (2018, Theorem 1.6)
@@
Combining Lemma A.3 and Lemma A.4 with $\|f\|_{W^{1,\infty}}\leq B$ yields for a constant $C>0$ for any $\tilde{\varepsilon}\in(0,1/2)$ that
@@
where $\tilde{\varepsilon}$ determines the approximation accuracy in Lemma A.4. For
@@
From Lemma A.2 we know that there is a ReLU network with no more than $1+\lceil C_{1}\log_{2}(\tilde{\varepsilon}^{-1})\rceil$ layers, $N^{d}+\lceil C_{2}(N+1)^{d}\log_{2}^{2}(\tilde{\varepsilon}^{-1})\rceil$ nonzero weights and $1+\lceil C_{3}(N+1)^{d}(\log_{2}^{2}(\tilde{\varepsilon}^{-1})\vee\log_{2}(\tilde{\varepsilon}^{-1}))\rceil$ neurons with the required properties. Inserting (23) and assuming $CB>\frac{1}{2}$ yields
@@
First we note that for $\Gamma>1,$ there is an $L>0,$ such that there is a $B>0$ with $12B<\Gamma-$ 1 and with $\hat{X}\sim\mathbb{P}^{*}$
$\sup_{W\in\operatorname{Lip}(L)}\mathbb{E}[W(\hat{X})-W(\hat{G}_{n}(Z))]=\sup_{W\in\operatorname{Lip}(L,2B)}\mathbb{E}[W(\hat{X})-W(\hat{G}_{n}(Z))].$
This $L>0$ exists as $[0,1]^{d}$ is bounded and adding a constant to any function $W\in\operatorname{Lip}(L)$ will not change the value of $\mathbb{E}[W(\hat{X})-W(\hat{G}_{n}(Z))].$
Then we get for every $G\in\mathcal{G}$ with the same reasoning as in the proof of Theorem 5.1